john likes 400 but not 300;he likes 100 but not 99;he likes 3600 but not 3700.
which does he like?

Answer by stanbon(75887)

John likes perfect squares:
4oo = 20^2
100 = 10^2
3600 = 60^2
Cheers,
stan H.